# DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents

### v2

@@ 4.2 Attack Transferability
Overall Transferability. Table 1 shows the transferability of the attacks under practical black-box settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. This is mainly because the adversarial text effectively reflects the attack goals. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
@@
DrunkAgent vs. Baseline Attacks. From Table 1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
4.3 Attack Stealthiness
Overall Stealthiness. Fig. 3 shows a overall distribution of the recommendation performance changes of the victim models before and after the attacks. The distribution includes all the victims and real-world datasets. To clearly demonstrate the differences, we adopt HR@3 to evaluate. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which signifies the attack does not disrupt the normal operation of RSs, making it difficult for users and platforms to detect and indicating its remarkably stealthy and unnoticeable.
@@ 4.3 Attack Stealthiness
Figure 3: Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
Attack Imperceptibility. Following the works (zhang2024stealthy, ; liu2024formalizing, ), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric. A lower perplexity indicates that the perturbed text remains close to natural language, thus appearing less suspicious. From Fig. 4, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only coherent and fluent but also semantically meaningful and stealthy.
@@
Figure 4: Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
4.4 Defense Strategies to Attacks
@@
Following the representative works (zhang2024stealthy, ; zhang2024agent, ), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses, as shown in Fig. 5. Moreover, from Fig. 5 and Table 1, DrunkAgent still performs better than the baselines after defenses were introduced. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words/phrases/sentences improves the semantically meaningfulness of adversarial text, making our attacks stronger.
5 Conclusion
@@
Figure 5: DrunkAgent’s Robustness to Defense Mechanisms.
@@
Appendix A Related Works
In recent years, data poisoning attacks (aka shilling attacks) [39, 18, 40, 41, 21, 19, 42, 12] have been fully investigated to perform robustness analysis on traditional RSs (e.g., matrix factorization [43], NCF [44], LightGCN [45]). Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items. However, they are less effective [13] even of limited applicable to the recent RSs that are empowered by LLMs [14]. The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations [15, 11], which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models. Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses [14]. Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models [13]. Meanwhile, jailbreak attacks [26, 27, 28] are proposed to undermine the general-purpose LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection. Nevertheless, such attacks have different attack goals from those item promotion attacks for RSs, making them hard to effectively transfer to the LLM-based RSs.
To fill the above gaps, two LLM-based RS attacks have been proposed [13, 14]. Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions. Zhang et al. [13] propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. This attack overcomes this limitation that classical text attacks [34, 35, 36, 37] fail on LLM-based RSs due to misaligned malicious objectives. Compared with those traditional shilling attacks, such an item representation attack is notably stealthy [13], as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews. However, its transferability is limited, due to the attack is not optimized on any black-box RSs [41, 17]. Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable. Another black-box attack CheatAgent [14] has been proposed. This work claims the prompt templates and user profiles on the victim models can be arbitrarily perturbed. However, such assumptions are impractical. The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible [13]; and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order [5]. Furthermore, CheatAgent is an untargeted attack, which is beyond the scope of this paper, due to we focus on targeted attacks.
Traditional RSs [46] are predominantly based on static models, where user preferences are learned from historical data in an offline manner and remain fixed during inference. These models do not adapt to evolving user behaviors or changing contexts after training. Vanilla LLM-based recommenders [15] follow a similar static paradigm: they generate recommendations through single-turn, stateless inference, relying on pre-defined prompt inputs to represent user preferences. In contrast, dynamic LLM-powered agentic recommenders [7, 11, 8] maintain internal states (e.g., memory mechanisms), allowing them to continuously adapt to and self-evolve over time based on sequential user interactions and simulated environmental feedback, even in offline experimental settings. This adaptivity enables personalized, context-aware, and temporally-sensitive recommendations. However, such dynamic mechanisms also introduce novel vulnerabilities that static models do not exhibit (e.g., memory corruption), which call for new robustness analyses tailored to the agent recommenders. While there are some attack methods on LLM-based agents [38], they often pursue goals that diverge from those relevant to RSs, leaving a gap in effective adversarial attacks for agent-driven recommenders.
To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs. As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-powered agentic RSs with the aim of promoting target items. We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles. The adversarial descriptions are optimized via a surrogate model to enhance the transferability and stealthiness of DrunkAgent.
Algorithm 1 Optimization Procedure of DrunkAgent
1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
[…]
4:     Quality Estimation Stage:
5:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
6:     Maintain top-$n$ candidates in descending order of scores
7:     Feature Integration Stage:
8:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
9:     Randomly exchange of the slices between pairwise texts
10:     Linguistic Enrichment Stage:
11:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
12:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
13:  end while
15:  Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
18:  end while
19:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
Appendix B Methodology
B.1 Supplements to Basic Adversarial Prompt Template
The main components of the prompt template are elaborately described in the following.
  1. 1.
    Style Definition $\widetilde{p_{s}}$. There are two common styles of prompts: task-oriented and role-oriented [47]. To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig. 2.
  2. 2.
    Task Goal $\widetilde{p_{g}}$. Since a recommender system is essentially a ranker [46], we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
  3. 3.
    Recommendation Instruction $\widetilde{p_{i}}$. We introduce highly detailed instructions to guide the model to perform accurate recommendations. For example, ‘Please sort CDs by considering the correlation between your preferences/dislikes and the features of the CDs. The higher the CD rank, the more you like it.’ in the shown example.
  4. 4.
[…]
B.2 DrunkAgent Algorithm
For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available. Importantly, to mitigate potential misuse risks, the released code excludes scripts for automatic attack execution and is provided solely for academic research under a non-commercial license. All the users are required to agree to responsible use guidelines before accessing the whole repository.
Appendix C Experiments
C.1 Supplements to Experimental Setup
Table 2: Statistics of Datasets
@@ C.1 Supplements to Experimental Setup
Table 3: Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
@@
C.1.1 Dataset Statistics
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work [5], we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions including 777 items, 99 users with 693 interactions including 619 items and 98 users with 588 interactions including 483 items from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy [44] is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
C.1.2 Baseline Methods
@@
  • Character-level Perturbations. DeepwordBug [34] and PuncAttack [35] manipulate texts by introducing typos and inserting punctuation, respectively.
@@
  • Word-level Perturbations. TextFooler [36] and BertAttack [37], respectively, aim to replace words with synonyms or contextually similar words.
@@
  • Sentence-level Perturbations. TrivialInsertion [13] inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. Here, $k=2$. ChatGPTAttack [13] increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs, where GPT-3.5-turbo is used as the original paper recommends and the maximum number of words is set to 30 here.
@@
We give some examples of adversarial descriptions of target items generated by the baselines in Table 3. Following the works [5, 7], the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig. 2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs [17, 48, 20], as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^3] https://github.com/CRIPAC-DIG/RecTextAttack [13] and PromptBench [^4] https://github.com/microsoft/promptbench [49].
C.1.3 Victim LLM-powered Agentic Recommender Systems
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs [11]. Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs [50, 7], AgentCF [5] emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering [44]. Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig. 1. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities. Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding. As can be seen on the right of Fig. 1, the recommendation agents conduct personalized recommendations based on the optimized memories. To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs. As such, AgentRAG and AgentSEQ are developed [5].
To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest [5], where gpt-3.5-turbo [^5] https://platform.openai.com/docs/overview is recommended to achieve the backbone LLMs. Specifically, such backbone delivers strong natural language understanding and generation capabilities that are sufficient for most recommendation-related tasks. In addition, it is cost-efficient, making it well-suited for multi-turn interactions where cost can quickly accumulate. Moreover, it has lower latency, supporting real-time responsiveness in interactive agent scenarios. To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents. Specifically, we introduce the Chain-of-Thought strategy [51] to carefully craft the prompt template $\mathcal{P}$. Given the context window constraints pointed out in [5], a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work. The design of recommendation agents is described below.
    $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (8)
    $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (9)
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2666https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
    $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$ (10)
@@ C.1.3 Victim LLM-powered Agentic Recommender Systems
Table 4: Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate Chain-of-Thought enhancement strategy.
@@
C.1.4 Configuration of DrunkAgent
As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for building the LLM backbone of the surrogate model $f_{\text{Agent}_{\Phi}}(\cdot)$. Its collection spans different sizes, from 8B, 70B to 405B. To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. OpenAI’s gpt-4-turbo [^8] https://platform.openai.com/docs/models/gpt-4-turbo [52] is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities. We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values.
C.1.5 Evaluation Metrics
We use two widely-used ranking metrics to evaluate the attack transferability and stealthiness: hit ratio (HR@$\mathcal{K}\uparrow$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}\uparrow$) [53, 19]. HR@$\mathcal{K}$ measures whether the target/test item is appeared in the top-$\mathcal{K}$ recommendation list of the normal users, while NDCG@$\mathcal{K}$ indicates the ranking position of the target/test item in the list. As such, a high HR@$\mathcal{K}$ and a NDCG@$\mathcal{K}$ demonstrate high attack effectiveness, and vice versa. To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3. We also introduce standard sentence perplexity score $\downarrow$ [13] that is a commonly-adopted metric to evaluate the attack imperceptibility. The perplexity measures a language model’s uncertainty by taking the exponential of the average negative log-likelihood of the words in a sequence, given their preceding context. Text with low perplexity is more fluent and natural, making it harder to detect as manipulated or adversarial, while text with high perplexity is more likely to be detected as unusual or abnormal.
C.2 Supplements to Performance Comparison
C.2.1 Attack Universality across Different Target Items.
Figure 6: Attack Universality across Different Target Items
Table 5: Ablation Studies of DrunkAgent
@@ C.2 Supplements to Performance Comparison
Table 6: Success Rates to Get the Target Agents Drunk
@@
C.2.2 Ablation Studies

### v3

@@ 4.2. Attack Transferability
Table 1 shows the transferability of the attacks under practical black-box settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. This is mainly because the three modules collaborate effectively, producing adversarial texts that accurately reflect the attack goals. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
4.2.2. DrunkAgent vs. Baseline Attacks.
@@
From Table 1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
Figure 3. Attack Universality across Target Items.
4.3. Attack Universality
@@
Figure 4. Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
@@
Figure 5. Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
4.4. Attack Stealthiness
4.4.1. Overall Stealthiness.
Fig. 4 shows a overall distribution of the recommendation performance changes of the victim models before and after the attacks. The distribution includes all the victims and real-world datasets. To clearly demonstrate the differences, we adopt HR@3 to evaluate. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which signifies the attack does not disrupt the normal operation of RSs, making it difficult for users and platforms to detect and indicating its remarkably stealthy and unnoticeable.
4.4.2. Attack Imperceptibility.
Following the works (Zhang et al., 2024c; Liu et al., 2024a), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric. A lower perplexity indicates that the perturbed text remains close to natural language and is more fluent, making it harder to detect as manipulated or adversarial. From Fig. 5, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only coherent and fluent but also semantically meaningful and stealthy.
4.5. Defense Strategies to Attacks
@@
Following the representative works (Zhang et al., 2024c, 2025), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses, as shown in Fig. 6. Moreover, from Fig. 6 and Table 1, DrunkAgent still performs better than the baselines after defenses were introduced. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words/phrases/sentences improves the semantically meaningfulness of adversarial text, making our attacks stronger.
@@
Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
5. Related Works
5.1. Data Poisoning Attacks on Traditional Recommender Systems
In recent years, data poisoning attacks (aka shilling attacks) (Li et al., 2016; Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023; Wang et al., 2024a) have been fully investigated to perform robustness analysis on traditional RSs (e.g., NCF (He et al., 2017), LightGCN (He et al., 2020)). Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items. However, they are less effective (Zhang et al., 2024c) even of limited applicable to the recent RSs that are empowered by LLMs (Ning et al., 2024). The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023; Huang et al., 2025a), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models. Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024). Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models (Zhang et al., 2024c).
5.2. Adversarial Attacks on LLM-based Recommender Systems
To fill the above gaps, two LLM-based RS attacks have been proposed (Zhang et al., 2024c; Ning et al., 2024). Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions. Zhang et al. (Zhang et al., 2024c) propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. This attack overcomes this limitation that classical text attacks (Gao et al., 2018; Formento et al., 2023; Jin et al., 2020; Li et al., 2020) fail on LLM-based RSs due to misaligned malicious objectives. Compared with those traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024c), as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews. However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024). Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable. Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed. This work assumes the prompt templates used by victim models can be arbitrarily perturbed, which is difficult to realize in practice, as such perturbations often require insider access or even system compromises. Such attacks aim to undermine the overall recommendation performance, making the attack perceptible (Zhang et al., 2024c). As such, CheatAgent is an untargeted attack, which is beyond the scope of this paper, since we focus on targeted attacks.
Unlike static traditional or vanilla LLM-based RSs, dynamic agentic RSs (Huang et al., 2024, 2025a, 2025b) maintain internal memory states that enable continuous adaptation to evolving user behaviors and changing contexts. This adaptivity yields more personalized and context-aware recommendations, but it also introduces new vulnerabilities that static models do not exhibit (e.g., memory corruption), calling for robustness analyses tailored to LLM-powered recommender agents.
To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs. As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-based agentic RSs with the aim of promoting target items. We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles. While inspired by static jailbreak attacks on general-purpose LLMs (Zou et al., 2023; Liu et al., 2024b, a), DrunkAgent establishes a new threat model and develops tailored optimization techniques for adapting to dynamic victim architectures and distinct attack goals in agentic recommendation scenarios.
6. Conclusion
@@
Appendix A Appendix
A.1. Supplements to Experimental Setup
Table 2. Statistics of Datasets
@@ A.1. Supplements to Experimental Setup
Table 3. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
@@
A.1.1. Dataset Statistics
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions including 777 items, 99 users with 693 interactions including 619 items and 98 users with 588 interactions including 483 items from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
A.1.2. Baseline Methods
@@
  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
@@
  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
@@
  • Sentence-level Perturbations. TrivialInsertion (Zhang et al., 2024c) inserts text content with several positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. ChatGPTAttack (Zhang et al., 2024c) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs.
@@
We give some examples of adversarial descriptions of target items generated by the baselines in Table 3. Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig. 2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^5] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024c) and PromptBench [^6] https://github.com/microsoft/promptbench (Zhu et al., 2023).
@@
Table 4. Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate Chain-of-Thought enhancement strategy.
@@
Table 5. Ablation Studies of DrunkAgent
@@
Table 6. Success Rates to Get the Target Agents Drunk
@@
A.1.3. Victim LLM-powered Agentic Recommender Systems
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs (Huang et al., 2025a). Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (He et al., 2017). Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig. 1. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities. Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding. As can be seen on the right of Fig. 1, the recommendation agents conduct personalized recommendations based on the optimized memories. To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs. As such, AgentRAG and AgentSEQ are developed (Zhang et al., 2024b).
To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest (Zhang et al., 2024b). To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents. Specifically, we introduce the Chain-of-Thought strategy (Wei et al., 2022) to carefully craft the prompt template $\mathcal{P}$. Given the context window constraints pointed out in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work. The design of recommendation agents is described below.
    (8) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    (9) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2777https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
    (10) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
A.2. Supplements to Performance Comparison
A.2.1. Ablation Studies